### v4

Abstract
Structured, or tabular, data is the most common format in data science. While deep learning models have proven formidable in learning from unstructured data such as images or speech, they are less accurate than simpler approaches when learning from tabular data. In contrast, modern tree-based Machine Learning (ML) models shine in extracting relevant information from structured data. An essential requirement in data science is to reduce model inference latency in cases where, for example, models are used in a closed loop with simulation to accelerate scientific discovery. However, the hardware acceleration community has mostly focused on deep neural networks and largely ignored other forms of machine learning. Previous work has described the use of an analog content addressable memory (CAM) component for efficiently mapping random forests. In this work, we develop an analog-digital architecture that implements a novel increased precision analog CAM and a programmable chip for inference of state-of-the-art tree-based ML models, such as XGBoost, CatBoost, and others. Thanks to hardware-aware training, X-TIME reaches state-of-the-art accuracy and 119$\times$ higher throughput at 9740$\times$ lower latency with >150×>150$\times$> 150 × improved energy efficiency compared with a state-of-the-art GPU for models with up to 4096 trees and depth of 8, with a 19W peak power consumption.
{IEEEkeywords}
Content Addressable Memories, Decision Trees, In-memory Computing, ReRAM
1 INTRODUCTION
[…]
Extracting relevant information from structured (i.e., tabular) data is the foundation of practical data science. Tabular data consists of a matrix of samples organized in rows, each of which has the same set of features in its columns, and is commonly used for medical, financial, scientific, and networking applications, enabling efficient search operations on large-scale datasets, making them one of the most common sources of data in the data science industry. Moreover, data collected in real-time for heterogenous sources, for example during scientific experiments, are organized in tabular data and with fast processing may lead to in-the-loop intelligent control.
While deep learning has shown impressive improvements in accuracy for unstructured data, such as image recognition or machine translation, tree-based machine learning (ML) models still outperform neural networks in classification and regression tasks on tabular data [1], including modern Transformers for tabular data [2]. This difference in model accuracy is due to robustness to uninformative features of tree-based models, bias towards a smooth solution of neural networks models, and in general the need for rotation variant models where the data is rotation variant itself, as in the case of tabular data [1]. With high model accuracy, ease of use, and explainability [3] tree-based ML models are greatly preferred by the data science community [4], with $>74\%$ of data scientists preferring to use these models while $<40\%$ choose neural networks in a recent survey.
However, tree-based models have garnered little attention in the architecture and custom accelerator research fields, particularly compared to deep learning. While small models are easily executed, performance issues for larger tree-based ML models arise due to limited inference latency both on CPU and GPUs [5, 6] originating from irregular memory accesses and thread synchronization issues. As large tree-based models ($>$1M nodes) are finding performance limits in existing hardware, they are demonstrating increasingly compelling uses in practical environments where high throughput and moderate latency are desired. For example, the winner of a recent IEEE-CIS Fraud detection competition used a tree-based ML model with 20M nodes [7]. Moreover, reducing inference latency is critical in scientific applications where for use in real-time processing, model latency must be $\sim$100 ns [8]. Similar low latency requirements for model inference arise in a real-time in-the-loop decision, data filtering systems, or streaming data scenarios, which are common across financial, medical, and cybersecurity domains [9, 7] where tabular data and tree-based ML also dominate. Thus, these models are reaching sizes and complexity levels for meaningful real-world tasks, as they are hitting limits on inference latency and throughput in current systems. Recently, custom accelerators for tree-based ML inference have been developed based either on FPGA [8, 9] or custom ASICs [10, 6]. A common architecture in most of the accelerators mentioned above implements look-up tables (LUTs) storing the attributes of a decision tree (DT) accessed multiple times during inference, with reads based on decision paths taken during the tree traversal. Along with low latency inference, an efficient hardware representation of trees is necessary to implement the required very large models.
One promising new architectural approach for improving tree-based model inference is in-memory computing, where data are stored and processed in the same location. This approach eliminates access, transfer, and processing of data in a separate compute unit [11], reducing latency. Previous research has shown that multiplication functions can efficiently be performed in crossbar architectures [12]. However, crossbar architectures are limited to matrix operations, which are not useful for inference on tree-based models. Another class of in-memory computing primitives is Content Addressable Memories (CAMs) [13, 14], high-speed and highly parallel circuits that search for a given input in a table of stored words and return the matched search input’s location. CAMs use relatively cheap peripherals, e.g. sense amplifier circuits rather than expensive ADCs required by crossbar arrays for in-memory computing operations. More recently, analog CAMs have been proposed [15] which leverage the analog states in resistive switching memories (RRAM) [16]. An analog CAM performs the comparisons of stored (analog) ranges with (analog) inputs, by computing if the input is within the range. Given that DTs are a set of conditional branches, analog CAMs naturally map DTs and perform highly parallel inference of decision forests [17, 18]. Our previous work showed the potential of this approach but was limited to component-level estimation, 4-bit precision, and random forest models [17], while the full architecture study in this work - which includes designs for higher precision Analog CAM operation, study of peripheral circuits and builds in support for multiple ML models convincingly demonstrates the feasibility of the initial concept.
This work presents X-TIME, an in-memory engine for accelerating tree-based ML models with analog CAMs. In this work we demonstrate the need for a vertical architecture design, 8-bit operation, and support for multiple tree-based ML models, overcoming the limitations of previous work [17]. The major contributions of this paper compared to previous work [17] are at the circuit level(1) a novel analog CAM cell design and operation to perform decision tree inference at doubled (i.e. 8-bit) resistive memory precision (i.e. 4-bit), (2) a complete design of the core and network operation, and at the architecture level (3) an end-to-end design targeting a large scale accelerator, (4) a compiler (X-TIME-C) supporting the mapping of multiple tree-based ML models to the HW starting from established intermediate representation [19], and at the co-design level (5) an HW-aware model training algorithm for reaching state of the art accuracy and (6) a complete simulation framework to benchmark and rigorous characterization comparison against NVIDIA V100 GPUs, FPGAs, and conventional digital accelerators. Our simulated evaluation shows up to 119$\times$ increased throughput at 9740$\times$ decreased latency with >150×>150$\times$> 150 × improved energy efficiency in a single 19W chip.
2 BACKGROUND
Figure 1: Schematic illustration of (a) a decision tree and (b) a decision tree ensemble.
2.1 Tree-based Machine Learning
A DT is a regression and classification ML model consisting of a set of nodes arranged in a tree-like structure. Each node can have multiple children nodes, but the most common case is a binary tree, where each node has only two children. The first node is usually called the root. Each node $j$ consists of a conditional branch where a single input feature element $f_{i}$ is compared with a trainable threshold $T_{j}$. The exception to this is the last node, or leaf, which stores the model’s prediction for a given input, and which can be a class (classification) or a scalar value (regression). A balanced tree can be defined by its depth $D$, which is equal to the maximum number of nodes in a root-to-leaf traversal. Fig. 1(a) shows an example of a DT with $L=4$ leaves and $D=2$.
Multiple DTs can be trained together as an ensemble for increased accuracy in classification and regression tasks. Fig. 1(b) shows a conceptual representation of a DT ensemble, where multiple trees are executed in parallel and their output combined with a reduction operation yields the model’s final result. Examples of DT ensembles model include Random Forests (RF) [20], eXtreme Gradient Boosting (XGBoost) [21], and Categorical Boosting (CatBoost) [22]. In the case of RF, the reduction operation consists of a majority voting, while in all the others leaf values of each DT are summed.
While training DT models is relatively fast when compared to, for example, a large deep neural network, the computations involved in inference are very irregular and not particularly suitable for CPUs and GPUs [6], becoming inefficient for very large models, with $>1$ M nodes [7]. While CPUs can run irregular control paths fairly well, their parallelism is limited. GPUs have a very large parallelism, but suffer from the irregularity of memory access patterns, as explained in detail in Supplementary Information LABEL:supp-si:section:gpu_inference.
Figure 2: Comparison of (a) a conventional CAM/TCAM and (b) an analog CAM
2.2 Analog Content Addressable Memory
CAMs are specialized memory devices that search a table of stored words (rows) or keys in parallel for a given input query. CAMs output a one-hot vector where a high value corresponds to a match between the input data query and stored row word key [13, 14]. In conventional CAMs, this output vector is converted into a single address by a priority encoder. Each CAM cell functions as both memory and bit comparison, and the input query is compared to each CAM word simultaneously to produce a high throughput, low latency few-cycle compare operation at the cost of high power and area. In addition to binary (0 and 1) values, ternary CAMs (TCAMs) have a third wildcard value, X (or don’t care), which can be used to compress the stored CAM table or perform partial match searches. When an X is stored or searched, that cell matches irrespective of whether the searched or stored value is 1 or 0. Fig. 2(a) shows a schematic illustration of a conventional CAM/TCAM circuit, where input queries are applied on the column Search Lines (SLs) and outputs are read along the row Match Lines (MALs). Essentially, an XNOR operation occurs between the query values and stored cell values in the corresponding column. The MAL $j_{th}$ row output is produced by a bit-wise AND operation between cells along the row and returns a match (high) or mismatch (low). The red boxes in Fig. 2(a) correspond to mismatches between input and stored value. In practice, MALs are initially pre-charged high and a mismatch in a cell activates a pull-down transistor connected to MAL to discharge it, such that a match is returned only if all pull-down transistors remain off during a search operation.
TCAMs are used ubiquitously in networking applications that require high-throughput and low-latency operations, for example for packet classification and forwarding, and access control lists [13]. While offering high performance, conventional SRAM-TCAMs are power and area-hungry and require 16 transistors. Recently, there has been significant interest in developing TCAMs with RRAM devices  [23, 24, 25] with compact size and low-power operation. As programming these devices is relatively slow, emerging TCAMs have targeted in-memory computing applications [26, 14] where stored value updates are infrequent.
Beyond direct digital memory operation, RRAM devices have also demonstrated ranges of stable analog states of $>16$ levels [27]. Leveraging this analog state capability, recent work proposed and demonstrated an analog CAM circuit design and operation [15] (conceptual representation in Fig 2(b)). In an analog CAM, ranges are stored in each cell, an analog input is applied vertically along columns and a match is returned along rows if the inputs are all within the stored range of each cell in a row. An analog CAM row implements
$\text{MAL}_{i}=\bigwedge_{j}{(T_{L,ij}\leq q_{j}<T_{H,ij})}\text{,}$ (1)
where $q$ is the query input vector applied to Data Line (DL), and $T_{L}$ and $T_{H}$ are the stored range thresholds for the CAM cell range lower and upper bounds, respectively. In the analog CAM circuit, no analog-to-digital conversion is required, eliminating an expensive component which is a limiting bottleneck of other analog in-memory computing primitives [28]. Multiple analog CAM circuits have been realized using different technologies such as RRAM devices [15] and ferroelectric memory devices [29]. In such circuits, two analog memory elements represent the lower and upper bound thresholds. Analog CAMs have been shown to map CAM tables more efficiently [15], as well as implement novel functions, such as computing distances for memory-augmented neural networks [29, 25], decision tree inference [17] and pattern learning [30].
Figure 3: DT example from Fig.1(a) with nodes mapped in an analog CAM and leaves mapped into a RAM.
2.3 Mapping tree models to analog CAMs
Fig. 3 shows the mapping of the decision tree in Fig. 1(a) to the Analog CAM. Each row represents a branch of the tree and each column corresponds to a different feature of the input vector $q$. Each node is mapped in the column corresponding to the feature used in the node’s comparison. The comparison’s threshold value is programmed in the analog CAM’s lower and/or upper bound, according to the direction, left or right, taken by the comparison’s result for the root-to-leaf path. A don’t care symbol is used if a feature is missing, by programming the full range (0 to 1 in our example) in the analog CAM cell. The feature vector (query) is applied on the DL and the output on the MAL is high for the matched branch, i.e. the traversed root-to-leaf path, in a single CAM operation. This mapping requires the number of CAM rows to match the number of tree leaves, and the number of CAM columns to match the number of features $N_{feat}$. MALs can be connected directly to the Word Lines (WLs) of a conventional SRAM storing leaf values to return the model’s prediction. Effectively, the analog CAM traverses root-to-leaf paths, and the RAM retrieves leaf values.
Previous results of this approach have shown significant speedups for RF inference [17] and reduced energy consumption compared to both GPU and state-of-art ASICs [31], although with limited architecture investigation for scaling and no flexibility to accommodate different dataset or model types. In this work, we present a complete architecture development and study for handling multiple DT ensemble algorithms such as RF, XGBoost, and CatBoost, and learning tasks such as regression, binary classification, and multi-class classification with the same architecture. In addition, we extend the use of the previously presented analog CAM [15, 17] circuit to support higher precision data by developing a flexible connectivity strategy and bit-arithmetic refactoring. We compare the performance of our architecture on state-of-the-art tabular dataset ML tasks against GPU, FPGA, and digital ASIC, showing a large improvement in latency and throughput while using a single chip with limited power consumption.
3 Hardware-aware model training
To ensure a targeted and relevant accelerator design, we explore the HW algorithm co-design space by assessing model types, sizes, and accuracy requirements in conjunction with different analog CAM HW and architecture parameters. First, we consider the limitations of the previously presented RF accelerator based on analog CAM [17]. The accelerator was limited to 4-bit precision, essentially due to the programming accuracy of ReRAMs, and it was benchmarked on a single dataset. However, different tree-based ML models would perform differently for different tasks, such as binary or multiclass classification or regression, and different datasets. The 4-bit operation might not be enough to capture more complex datasets that would require a larger number of bins for each feature. Finally, the maximum depth of each tree and the number of trees, corresponding to the number of words in the analog CAM and the number of analog CAM arrays, should be optimized on a variety of datasets to ensure a general well-performing accelerator. Thus we performed an in-depth characterization of tree-based ML models, namely RF, XGBoost, and CatBoost, performance on several datasets to understand (1) the necessity for mapping diverse models to the accelerator, (2) the size of each analog CAM macro and total number of macros in the accelerator and (3) the required bit precision for the analog CAM operation.
Table 1: Datasets and models characterization
Dataset ID Task Samples $\boldsymbol{N_{feat}}$ $\boldsymbol{N_{classes}}$ Model $\boldsymbol{N_{trees}}$ $\boldsymbol{N_{leaves,max}}$
Churn modelling[32] 1 Binary Classification 10000 10 2 CatBoost 404 256
Eye movements[33] 2 Multiclass Classification 10936 26 3 XGBoost 2352 256
Forest cover type[34] 3 Multiclass Classification 581012 54 7 XGBoost 1351 231
Gas concentration[35] 4 Multiclass Classification 13910 129 6 Random Forest 1356 217
Gesture phase segmentation[36] 5 Multiclass Classification 9873 32 5 XGBoost 1895 256
Telco customer churn[37] 6 Binary Classification 7032 19 2 XGBoost 159 4
Rossmann stores sales[38] 7 Regression 610253 29 N/A XGBoost 2017 256
We perform a survey on the size of typical tabular datasets [1], finding that usually $N_{feat}<100$ and often feature pre-processing reduces the number of $N_{feat}$ significantly before applying inputs to the decision tree ensemble. Nevertheless, we include an outlier dataset with $N_{feat}=130$[35] to show how to handle larger input vector sizes with input vector segmentation, as will be described later. Table 1 shows a characterization of the selected datasets.
Figure 4: (a) RMSE as a function of model size in the case of Rossmann stores sales[38] for multiple configurations.(b) Best accuracy of each dataset for multiple configurations.
First, we trained all the models optimizing the hyperparameters without limitations, similar to what would happen if training the model for the GPU. Then, we studied the impact of bit precision. Given that from implementing 4-bit to 8-bit operation the analog CAM area would roughly double, as explained in Section 4.2, we doubled the number of available trees in that case. Finally, we limited the training to RF to demonstrate that it underperforms compared to other models, such as XGBoost, thus a general architecture that accommodates gradient-boosted models, as opposed to the previously realized accelerator [17] for RF-only is vital.
Fig. 4(a) shows as an example the Root Mean Square Error (RMSE) as a function of the model size expressed as the size of an analog CAM-based hardware, namely $N_{trees}\times 2^{D}$, with $D$ maximum depth and $N_{trees}$ number of trees in the case of Rossmann stores sales dataset[38]. We repeated the same experiments for all datasets (Fig. LABEL:supp-si:fig:dse and LABEL:supp-si:fig:dse2), picking as design choices for our architecture $N_{trees}=4096$ and $D=8$, corresponding to a total of 4096 cores of with $2562^{8}=$ 256 words each. The dashed line corresponds to our selected maximum model size, and thus hardware size and light blue area correspond to the model that we can map to the hardware. It is possible to see how the accuracy saturates at a certain model size regardless of the configuration chosen (i.e. 4 or 8-bit, RF-only, or any tree-based model).
Results show that RF models are insufficient to reach good precision, and 8-bit inference can increase accuracy up to 2% for classification and a $\sim$20reduction on RMSE for regression. Fig. 4(b) shows the accuracy and RMSE for multiple datasets assuming constraint-free hyperparameters optimization (GPU), limited to our design choice with 8-bit (X-TIME 8-bit) and 4-bit (X-TIME 4-bit), and using only RF models (RF Only) demonstrating that (1) 8-bit can match the constraint-free design, (2) 4-bit is not enough for some datasets and (3) RF is underperforming significantly the best models for some datasets. A printout of the data can be also found in Table  LABEL:supp-si:tab:accuracy.
The final model and parameter selections for each dataset are reported in Table 1. Given that the hardware cost scales aggressively with $D$ but linearly with $N_{trees}$ (Fig. LABEL:supp-si:fig:cost), and the accuracy saturates for $D>8$, we designed X-TIME for being able to perform inference of models up to $N_{trees}=4096$, $D=8$, and $N_{bit}=8$. If more trees are needed, multiple X-TIME chips can be operated in parallel, similarly to how multiple cores are operated in parallel.
4 X-TIME Architecture
Figure 5: Architecture hierarchy
Fig. 5 shows the internal architecture of an X-TIME single-chip accelerator. The accelerator is composed of a group of cores connected by an H-tree NoC topology converging to a co-processor (CP). Each core contains an analog CAM array acting as a processing element (PE-CAM) and an SRAM memory. To map models larger than a single array, each PE-CAM has stacked arrays (extended row-wise) and queued arrays (extended column-wise) to handle more and larger words, respectively. Stacked arrays share the same peripherals, namely the Pre-Charger (P-Ch), the Digital Analog Converter (DAC), the Sense Amplifier (SA), and the Match-Line Register (ML-REG). The connection of queued arrays ensures that only previously matched lines are charged by making the ML-REG of array $i$ feed the P-Ch of the array $i+1$.
4.1 Inference
The X-TIME Compiler (X-TIME-C) maps the trained trees to the cores according to Fig. 3. X-TIME-C is based on Treelite [19], a universal model exchange and serialization format for tree-based models. Treelite outputs a C-like code with conditional statements representing each tree in the ensemble as an intermediate representation for further compilation to the hardware. X-TIME-C converts this code for analog CAM usage, obtaining a table of size $L\times(2N_{feat}+3)$ with each row storing the lower/upper bound for each feature, the leaf value, class ID/label (in the case of multi-class classification) and tree ID, representing the model ensemble for our architecture. X-TIME-C assigns trees to cores via round-robin, directly mapping the first $L\times(2N_{feat})$ columns of a tree to the lower and upper bound of the analog CAM circuit, and the third-last column to the SRAM. Class and tree ID are uniquely represented in the core address. Fig. LABEL:supp-si:fig:compiler shows a schematic visualization of X-TIME-C compilation flow.
The largest ensemble that can be mapped to X-TIME is constrained by the maximum number of leaves $N_{leaves,max}$ in all trees and the number of available cores $N_{C}$.
During inference, a feature vector is sent through the network to a core and searched in the PE-CAM. Each queued analog CAM block receives a different portion of the feature vector which is converted into analog voltage values by a DAC. A logical AND operation is inherently performed between the queued arrays, similar to ML segmentation used in modern TCAMs to maintain high throughput where MLs are divided into sections of shorter words with an AND between individual ML results. Connecting multiple TCAM arrays with common MLs is logically equivalent to a larger TCAM, with the bit-wise AND operation between TCAM cells on the same ML now extended between arrays. In our case, the analog CAM stores a branch split into multiple analog CAM arrays and reconstructed using the logical AND, such that only if all branch nodes arranged in multiple arrays are matched will a match be returned on the corresponding MAL.
The final output of the PE-CAM is stored in a register and used to address the SRAM, storing the leaf value of each branch.
Leaf values coming from multiple cores are summed with a programmable adder tree to perform the reduction operation. Finally, the resulting leaf value is propagated to the CP which performs a global reduction and computes the model prediction. The latter is usually a simple operation, such as a majority voting or a threshold comparison. In the case of a model larger than what can be mapped in a single X-TIME chip, we foresee the possibility of multiple CPs in multiple chips orchestrating a reduction similar to the adder tree in the chip network.
X-TIME is an inference accelerator, but models in production might be retrained and updated, due to data and/or concept drift, up to every 100 days [39]. While ReRAM endurance might be limited to 105-108 cycles depending on the technology [40] given the large retraining cycle, the lifetime of the accelerator is significantly larger than the lifetime of typical systems, making ReRAM endurance not a strong requirement.
4.2 Analog CAM with improved precision
Figure 6: (a) Circuit schematic and (b) block diagram of a AND-type 4-bit 6T-2M analog CAM [15]. (c) Block diagram of a 4-bit OR-type analog CAM. (d) Block diagram of an 8-bit analog CAM.
As a result of experiments in section 3 we found out that typically 8 bits precision, corresponding to 256 bins for each feature, is required for matching floating point (FP) accuracy inference of tree-based ML models.
While analog CAM circuits offer the possibility of mapping arbitrary ranges, RRAM device reliability poses a challenge in supporting more than 16 levels or 4 bits [17]. In the case of crossbar arrays mapping algebraic matrices, such as a neural network layer, multiple techniques have been proposed for overcoming limited device precision. For example, with bit slicing [12] by using two $M$ bits RRAMs, $N=2^{2M}$ bits can be represented by connecting two devices on the same row (Fig. LABEL:supp-si:fig:slicing(a)). In crossbar arrays, it is also straightforward to arbitrarily adjust the input precision by applying binary pulses and performing SHIFT and ADD operations at the output [12].
Unfortunately, such techniques do not apply to analog CAM arrays. The analog CAM encodes a unary representation and by connecting two analog CAM cells on the same MAL to represent one macro-cell, it is only possible to double the number of levels (i.e. $2^{N_{bits}}$) rather than the number of bits. Thus given RRAMs with $M=4$ bits, 16 analog CAM cells are needed for representing $N=8$ bits, leading to an exponential overhead (Fig. LABEL:supp-si:fig:slicing(b)). Here, we propose a novel solution to increase the precision with the same base analog CAM cell circuit, which requires 2 analog CAM cells and 2 clock cycles for performing 8-bit search operation with $M=4$ bits RRAMs.
To demonstrate, let us first take the case where the lower bound threshold $T_{L}$ is needed to be mapped and operated with $N=8$ bit of precision but we only have $M=4$ bits of precision in our analog CAM and RRAMs devices. The lower bound threshold can be sub-divided into its most significant bit representation $T_{LMSB}$ and least significant bit representation $T_{LLSB}$, each one made of $M$ bits and written as $T_{L}=2^{M}T_{LMSB}+T_{LLSB}=16T_{LMSB}+T_{LLSB}$. Each analog CAM cell computes the conditional operation $\textrm{MAL}=q\geq T_{L}$, where $q$ is the input query applied to the DL. We similarly divide the input $q$ into most and least significant representations $q_{MSB}$ and $q_{LSB}$, such that $q=2^{M}q_{MSB}+q_{LSB}=16q_{MSB}+q_{LSB}$ and we can obtain the result of MAL by computing
$\begin{split}[(q_{MSB}\geq T_{LMSB})\wedge(q_{LSB}\geq T_{LLSB})]\\ \vee(q_{MSB}\geq T_{LMSB}+1)\end{split}$ (2)
or alternatively, rearranging the terms:
$\begin{split}[(q_{MSB}\geq T_{LMSB}+1)\vee(q_{LSB}\geq T_{LLSB})]\\ \wedge(q_{MSB}\geq T_{LMSB}).\end{split}$ (3)
Note that Eq. 3 is more CAM-friendly since it is a conjunction of two terms that can be appended to the same MAL. The same calculation derives the upper bound $T_{H}$ conditions and with an AND operation between this and equation (3), the full computation $\textrm{MAL}=T_{L}\leq q<T_{H}$ results in:
$\begin{split}[(q_{MSB}\geq T_{LMSB}+1)\vee(q_{LSB}\geq T_{LLSB})]\\ \wedge(q_{MSB}\geq T_{LMSB})\\ \wedge[(q_{MSB}<T_{HMSB})\vee(q_{LSB}<T_{HLSB})]\\ \wedge(q_{MSB}<T_{HMSB}+1).\end{split}$ (4)
Analog CAMs cells arranged on the same MAL perform a bit-wise AND ($\wedge$) operation between them, thus a circuit addition is needed for performing the OR ($\vee$) operation. We start from the conventional analog CAM cell shown in Fig. 6 (a). The lower bound (yellow, T1, $M_{l}o$) is encoded in a 1-transistor-1-resistor structure. If the input is high enough, it turns on T1, lowering the gate node of the pulldown T2 which stays off, resulting in a match. If the input voltage is low or alternative $M_{l}o$ conductance is high, T2 is turned on. Thus the lower bound checks if $SL_{1}>f(M_{l}o)$. A similar discussion can be applied to the higher bound (orange, T3, $M_{h}i$, T4, T5), in which an inverter (T4,T5) is added between the input voltage divider (T3,$M_{h}i$) and the pulldown (T6) to represent the higher bound and checking if $SL_{2}\leq(M_{h}i)$. Fig. 6 (b) shows a more compact circuit schematic for the analog CAM in Fig. 6(a), with comparators driving the pulldown representing the 1T-1M and 1T-1M-1INV structures. We refer to this circuit as AND-type, given the AND operation between the lower and upper bound. Similarly, an OR-type circuit can be realized (Fig. 6 (c)) [30], effectively performing an out of range matching. Finally, by combining the two structures, it is possible to realize the general circuit of Fig. 6 (d) which can implement Eq. 4, by applying opportunely inputs to the comparator (see Supplementary Information LABEL:supp-si:section:8bit-acam) Note that separate inputs for each cell are required for the programming operation, thus no extra wiring to the cell is required.
5 Benchmark Results
5.1 Comparison with state-of-the-art
Figure 7: Comparison of (a) latency and (b) throughput of this work, GPU, FPGA, and fully digital ASIC accelerator from Fig. LABEL:supp-si:fig:digital_accelerator.
We benchmarked X-TIME using our open-sourced cycle-approximate simulator [^1] https://github.com/HewlettPackard/X-TIME on a wide range of models and dataset types (Table 1), and compared latency and throughput (see Fig. 7(a,b)) to GPU, specifically NVIDIA V100, FPGA [9] and a digital accelerator based on small SRAMs and similar to Booster [6] (Fig. LABEL:supp-si:fig:digital_accelerator). More details on the methodology can be found in Supplementary Information LABEL:supp-si:section:soa, LABEL:supp-si:section:methods and Table LABEL:supp-si:tab:benchmark. When possible, we applied input batching and replicated trees to increase throughput (Fig. LABEL:supp-si:fig:batching). As shown, our architecture produces drastically reduced latency, frequently to $\sim$100ns compared to GPU at 10 $\mu$s to ms, while also demonstrating significantly improved throughput by 10-120$\times$ across most datasets. Our architecture shines for large binary classification or regression models with a peak improvement in latency by 9740×9740$\times$9740 × at 119$\times$ higher throughput for the Churn modeling dataset inference [32] compared to GPU. As tree search is parallel in each analog CAM array of each core, as well as the in-network reduction, latency, and throughput are constant with $N_{trees}$ and $N_{leaves}$ with a delay solely dependant on $N_{feat}$ and $N_{classes}$. In contrast, on GPU there is a linear dependence on $N_{trees}$ and $D$, which limits the performance for large models. In the FPGA implementation, namely FAXID, there is a linear dependence on $N_{trees}$. The SRAM-based accelerator has a linear dependence on $D$ (since multiple memory accesses are required) and $N_{trees}$ when more than a tree needs to be mapped in a core.
Figure 8: Throughput as a function of number of trees $N_{trees}$ and depth $D$ for various accelerators. Note that in the case of X-TIME and the FPGA implementation, there is no dependence on the depth, removing the load imbalance issue, while only X-TIME shows no dependence on the number of trees.
As an example, we studied the throughput as a function of $N_{trees}$ and $D$ for GPU, FPGA, SRAM-based accelerator, and X-TIME as shown in Fig. 8, for a multiclass classification dataset [34], demonstrating the predicted trend. We note that the SRAM-based accelerator still suffers from load imbalance between trees, similar to GPU, in which trees with different depths have different inference times, which results in the need for synchronization before the reduction operation. As an energy comparison with GPU, we considered a worst-case scenario. Fig. LABEL:supp-si:fig:energy shows the energy saving in comparison by considering the peak power consumption (19 W) and idle power consumption (25 W) in the case of X-TIME and GPU, respectively. Even by assuming a full chip utilization for X-TIME and no energy consumption for the tree inference on GPU, X-TIME is up to 157$\times$ more energy efficient.
5.2 Sensitivity to noise
Figure 9: Accuracy and RMSE for different noise configurations
Analog hardware is subject to noise and non-idealities. However, tree-based ML ensembles are resilient to variation thanks to the averaging mechanism happening between trees of multiple boosting rounds. Fig. 9 shows the average accuracy of the models for different noise profiles, namely no noise (ideal), only ReRAM or DAC noise, and both noise sources, over 100 inference experiments. A printout of the data can also be found in Table  LABEL:supp-si:tab:noise. We characterized the noise of TaOx-based ReRAM devices at various conductance values ranging from $<0.1$ nS to 500 $\mu$S [27], by reading them 1000 times and extracting mean and standard deviation. We measured a maximum relative standard deviation of noise of $\sigma_{G}/G=0.1$. While DAC designs can reach higher precision, only 4-bits are needed for X-TIME inference. Nevertheless, we considered a pessimistic standard deviation of noise for the DAC $\sigma_{DAC}=50mV$. Results indicate no perturbation in the accuracy and RMSE when either ReRAM or DAC noise is applied, thanks to our hardware-aware training and inherent averaging mechanisms of ensemble models. In the case of regression and with both noise sources applied there is an RMSE perturbation, but still reduced compared to the 4-bit operation (Fig. 4).
Intuitively, with the thresholding mechanism of Analog CAM, it is possible to program conductances with a separation such that the probability of exceeding the threshold is relatively low, leading to good noise resistance performance.
6 Conclusion
In this work, we presented a novel architecture for accelerated tree-based ML inference with in-memory computing based on a novel increased precision analog non-volatile CAMs. Analog CAMs enable novel applications of in-memory computing beyond matrix multiplies. Moreover, the required peripherals are significantly reduced compared to crossbar-based computing primitives. Our approach maps decision trees to analog CAM ranges, which directly compute the matched results in parallel, given an input feature vector. Leaf values are stored in a local SRAM and accumulated at the core level when multiple trees are mapped in the same core. Outside the core, an H-tree NoC connects 4096 cores and a co-processor in a single chip. An in-network computing structure accumulates leaf values as data flows to limit data movement. The on-chip network is re-configurable to map different models (regression, multi-class classification, etc) and inference techniques, such as input batching. An HW-aware training routine ensures to gets state of the art accuracy by leveraging quantization and optimizing the number of trees and maximum depth. Our results show significant improvement in latency – by almost four orders of magnitude, compared with an NVIDIA V100 GPU, at a throughput up to 120$\sim$ 120$\times$ 120 × higher across a wide range of dataset types and model topologies. Further, the worst case scenario showed up to 150×150$\times$150 × improvement in energy efficiency compared to GPUs with our approach.